OpenStudy (anonymous):

A=400(1+.007/12)12*8

OpenStudy (anonymous):

The result is 3222.4

OpenStudy (shadowfiend):

So 0.007 is 7/1000. 7/1000/12 is 7 / 12000. Since multiplication is associative, you can also reorder the multiplication a bit, so we do all the whole numbers first: 400*12*8*(1+7/12000) = 38400*(1+7/12000) Since 1 is the same as 12000/12000, we can make an improper fraction: 38400 * (12007/12000) = 38422.4